OpenStudy (anonymous):

I need help!! please!! Question:: Find the height of a cone with volume 66.46 cm3 and radius 2.75 cm

OpenStudy (anonymous):

\[ V = \frac 1 3 \pi r^2 h \]

OpenStudy (anonymous):

\[ h = \frac { 3 V}{\pi r^2} \]

OpenStudy (anonymous):

\[h=\frac { 3 (66.46)}{3.14 (2.75^2)} \]

OpenStudy (anonymous):

\[h=8.39627 \]

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

yw